1.) Given P(B) = 0.27, P(A and B) = 0.15, P(A or B) = 0.47, what is P(A)?

Answer in decimal form. Round to 2 decimal places as needed.

2.)Given that P(A) = 0.47, P(B) = 0.07, and P(A and B) = 0.0329, are events A and B independent?
Answer #1

Solution:

1)

P(A or B ) = P(A) + P(B) - P(A and B )

P(A) = P(A or B ) - P(B) + P(A and B )

= 0.47 - 0.27+ 0.15

= 0.35

2)

P(A and B ) = P(A) * P(B)

= 0.47 * 0.07

= 0.0329

Therefore, the events A and B independent events.